OpenStudy (anonymous):

Charles can do a job in 1 hour, Bill can do the same job in 2 hours, and Bob can do the job in 3 hours. How long does it take them to do the job working together?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

Always express how much of the work each person can do in 1 unit of time (hours in this case) Charles does 1/1 part of the work, Bill 1/2 of the work, Bob 1/3 of the work in 1 hour so together they finsih (1+1/2+1/3) work in 1 hour..now just inverse the fraction to get the no. of hours (less than 1 here) in which they complete the work together

OpenStudy (anonymous):

could you help me out with this problem?

OpenStudy (anonymous):

yeah so they complete (1+1/2 +1/3) of the work in 1 hour..so work done in 1 hour= (6+3+2)/6=11/6 of the work..or they finish the work together in 6/11 of an hour or ~32.7 mins.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

now I understand

OpenStudy (anonymous):

next problem?

OpenStudy (anonymous):

I have a question to victor.ray's answer. How did you get the (6+3+2)/6

OpenStudy (anonymous):

(1/1 +1/2 + 1/3)= (6/6 + 3/6 + 2/6)= (6+3+2)/ 6-->The method for adding fractions with unequal denominators is to convert the denominators of each to the LCM (Lowest common multiple) of all denominators by multiplying both above and below with the required number